OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in simplifying : 3x^2-3x-60/x^2+9x+20. Be sure to list restrictions. [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (anonymous):

factoring out the 3 from the numerator leaves you with 3[x^2-x-20], then factoring both numerator and denominator you end up with\[\frac{ 3[(x-5)(x+4) }{ (x+4)(x+5) }\] as you can see if x=-4 or -5 denominator will be zero, and you can't divide by zero, those are you restrictions. you can also simplify: \[\frac{ 3(x-5)}{(x-5) }\]

OpenStudy (anonymous):

sorry denominator should be (x+5)

OpenStudy (anonymous):

@arjont thank you so much !